# MatchFixAgent: Language-Agnostic Autonomous Repository-Level Code Translation Validation and Repair

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/2018/06
@@ 3.2.3. IO Analysis
The IO Analyzer is prompted to evaluate whether the observable input-output behavior of the source and target fragments is semantically equivalent. The prompt includes an IO equivalence definition, which assess five dimensions: (1) semantic equivalence of accepted inputs, (2) consistency of produced outputs, (3) preservation of side effects (e.g., file operations, network calls, or global state modifications), (4) uniform handling of edge cases, and (5) similarity in performance-critical complexity. The LLM is prompted also prompted to produce a plausible input that would trigger dissimilar IO behavior if it believes the translation is inequivalent. This methodology catches inequivalences such as differing error messages, inconsistent encoding assumptions, or missing side effects—often overlooked by structural analyses (§3.2.13.2.2) alone.
@@ 4.1. Experimental Setup
AlphaTrans (Ibrahimzada et al., 2025) cli (Foundation, 2025b) Java Python 273 37841 100 372 201
csv (Foundation, 2025c) 235 33072 100 392 278
fileupload (Foundation, 2025d) 192 3567 100 246 185
validator (Foundation, 2025e) 646 41605 95.5 216 164
@@
colorsys (Team, 2025l) 8 120 96.3 67900 32300
heapq (Team, 2025n) 22 189 78.1 67900 32300
html (Team, 2025o) 44 684 65.3 67900 32300
mathgen (Weiler, 2025) 81 735 93.8 711 183
rbt (Algorithms, 2025b) 27 366 88.1 203000 47000
strsim (Luo, 2025) 64 654 20.8 1014 125
toml (Pearson, 2025) 72 1206 62.5 1126 192
@@
We evaluate MatchFixAgent on benchmarks used in prior work on automated repository-level translation. Each benchmark problem is a translation pair: the source function and the corresponding translation. The task for each benchmark problem is to give a verdict on the functional equivalence of pairs in two different PLs, and repair translation in case of equivalence. Our subjects are open-source repository-level translations with equivalence verdicts available [^4] We exclude rule-based transpilers as they do not propose a validation mechanism, i.e., their translation is (theoretically) correct by construction. from the peer-reviewed literature [^5] This criterion excludes techniques such as RustMap (Cai et al., 2025) and C2SaferRust (Nitin et al., 2025).. We make selections across a diverse set of PL pairs.
@@ 4.5. RQ4: Ablation Study
We perform another study by removing only the semantic analysis results when prompting the test generator and verdict agents. We use the same set with 1,0911{,}091 instances from the previous ablation to perform this study. Figure 8 illustrates the result of our second ablation. The results indicate that the performance of MatchFixAgent significantly drops by $39.7\%$ without the six semantic analyses. Furthermore, we observe that the test generator agent without semantic analyzer is more costly and on average spends $3.9\%$ ($66.3K$ instead of $63.7K$), $6.2\%$ ($136K$ instead of $128K$), $4.2\%$ ($9.4M$ instead of $9.0M$), and $7.5\%$ ($61.28h$ instead of $56.71h$) more turns/interactions, input tokens, output tokens, and time, respectively. This is expected since the purpose of semantic analysis is to provide the necessary context for the test generator agent, without the need for the agent to explore and figure it out for itself. Since each semantic analysis is a single call to the LLM, it is much cheaper than the agents that iteratively analyze the codebase.
@@ 5. Related Work
Translation Validation and Repair. Existing automated translation validation techniques either rely on test execution (Pan et al., 2024; Zhang et al., 2025; Ibrahimzada et al., 2025; Shetty et al., 2024; Ou et al., 2024; Ziftci et al., 2025; Xue et al., 2025; Yang et al., 2024b), formal methods (Yang et al., 2024c; Nitin et al., 2024), or fuzzing (Eniser et al., 2024) for translation validation. Abid et al. (Abid et al., 2024) and AlphaTrans (Ibrahimzada et al., 2025) leverage GraalVM (Oracle, 2025) and language interoperability to execute code in both source and target PL for translation validation. Other tools like Oxidizer (Zhang et al., 2025) and Syzygy (Shetty et al., 2024) instrument programs to extract input-output pairs and use in target PL for validation. Skel (Wang et al., 2025a) validates translations through test execution by converting Python tests to JavaScript simply through string replacement. This suffices since source Python tests are simple function calls and value comparisons. For automated translation repair, most tools (Zhang et al., 2025; Wang et al., 2025a; Ibrahimzada et al., 2025; Shetty et al., 2024; Ou et al., 2024; Yang et al., 2024b; Pan et al., 2024) rely on simple reprompting of LLMs with execution feedback, which has proven ineffective. Specifically, AlphaTrans (Ibrahimzada et al., 2025) performs independent reprompting of suspicious fragments based on execution trace, while Skel (Wang et al., 2025a) requires a user for manually repairing translation bugs.

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Foundations of Software Engineering; July 05–09, 2026; Montreal, Canadaisbn: 978-1-4503-XXXX-X/2018/06
@@ 3.2.3. IO Analysis
The IO Analyzer is prompted to evaluate whether the observable input-output behavior of the source and target fragments is semantically equivalent. The prompt includes an IO equivalence definition, which assess five dimensions: (1) semantic equivalence of accepted inputs, (2) consistency of produced outputs, (3) preservation of side effects (e.g., file operations, network calls, or global state modifications), (4) uniform handling of edge cases, and (5) similarity in performance-critical complexity. The LLM is prompted also prompted to produce a plausible input that would trigger dissimilar IO behavior if it believes the translation is inequivalent. This methodology catches inequivalences such as differing error messages, inconsistent encoding assumptions, or missing side effects—often overlooked by structural analyses (§3.2.1, §3.2.2) alone.
@@ 4.1. Experimental Setup
AlphaTrans (Ibrahimzada et al., 2025) cli (Foundation, 2025b) Java Python 273 37841 93.8 372 201
csv (Foundation, 2025c) 235 33072 88.1 392 278
fileupload (Foundation, 2025d) 192 3567 20.8 246 185
validator (Foundation, 2025e) 646 41605 62.5 216 164
@@
colorsys (Team, 2025l) 8 120 100 67900 32300
heapq (Team, 2025n) 22 189 100 67900 32300
html (Team, 2025o) 44 684 95.5 67900 32300
mathgen (Weiler, 2025) 81 735 100 711 183
rbt (Algorithms, 2025b) 27 366 96.3 203000 47000
strsim (Luo, 2025) 64 654 78.1 1014 125
toml (Pearson, 2025) 72 1206 65.3 1126 192
@@
We evaluate MatchFixAgent on benchmarks used in prior work on automated repository-level translation. Each benchmark problem is a translation pair: the source function and the corresponding translation. The task for each benchmark problem is to give a verdict on the functional equivalence of pairs in two different PLs, and repair translation in case of equivalence. Our subjects are open-source repository-level translations with equivalence verdicts available [^4] We exclude rule-based transpilers as they do not include a validation mechanism, i.e., their translation is (theoretically) correct by construction. from the peer-reviewed literature [^5] This criterion excludes techniques such as RustMap (Cai et al., 2025) and C2SaferRust (Nitin et al., 2025).. We make selections across a diverse set of PL pairs.
@@ 4.5. RQ4: Ablation Study
We perform another study by removing only the semantic analysis results when prompting the test generator and verdict agents. We use the same set with 1,0911{,}091 instances from the previous ablation to perform this study. Figure 8 illustrates the result of our second ablation. The results indicate that the performance of MatchFixAgent significantly drops by $39.7\%$ without the six semantic analyses. Furthermore, we observe that the test generator agent without semantic analyzer is more costly and on average spends $3.9\%$ ($66.3K$ instead of $63.7K$), $6.2\%$ ($136K$ instead of $128K$), $4.2\%$ ($9.4M$ instead of $9.0M$), and $7.5\%$ ($61.28\,\mathrm{h}$ instead of $56.71\,\mathrm{h}$) more turns/interactions, input tokens, output tokens, and time, respectively. This is expected since the purpose of semantic analysis is to provide the necessary context for the test generator agent, without the need for the agent to explore and figure it out for itself. Since each semantic analysis is a single call to the LLM, it is much cheaper than the agents that iteratively analyze the codebase.
@@ 5. Related Work
Translation Validation and Repair. Existing automated translation validation techniques either rely on test execution (Pan et al., 2024; Zhang et al., 2025; Ibrahimzada et al., 2025; Shetty et al., 2024; Ou et al., 2024; Ziftci et al., 2025; Xue et al., 2025; Yang et al., 2024b; Ke et al., 2025; Dehghan et al., 2025), formal methods (Yang et al., 2024c; Nitin et al., 2024), or fuzzing (Eniser et al., 2024) for translation validation. Abid et al. (Abid et al., 2024) and AlphaTrans (Ibrahimzada et al., 2025) leverage GraalVM (Oracle, 2025) and language interoperability to execute code in both source and target PL for translation validation. Other tools like Oxidizer (Zhang et al., 2025) and Syzygy (Shetty et al., 2024) instrument programs to extract input-output pairs and use in target PL for validation. Skel (Wang et al., 2025a) validates translations through test execution by converting Python tests to JavaScript simply through string replacement. This suffices since source Python tests are simple function calls and value comparisons. For automated translation repair, most tools (Zhang et al., 2025; Wang et al., 2025a; Ibrahimzada et al., 2025; Shetty et al., 2024; Ou et al., 2024; Yang et al., 2024b; Pan et al., 2024) rely on simple reprompting of LLMs with execution feedback, which has proven ineffective. Specifically, AlphaTrans (Ibrahimzada et al., 2025) performs independent reprompting of suspicious fragments based on execution trace, while Skel (Wang et al., 2025a) requires a user for manually repairing translation bugs.